# Shot-noise in strongly-correlated double quantum spin Hall edges

### v2

@@ Abstract
We consider the effects of interactions on the edges of “double” quantum spin Hall insulators (DQSHIs), motivated by recent experiments on moiré twisted metal dichalcogenides. Without interactions, a DQSHI can be understood as two identical copies of a conventional quantum spin Hall insulator. If interactions are present and $s^{z}$-spin is conserved, we show that there are two possible phases for the DQSHI edge. First is a weakly-correlated edge which has two pairs of helical modes and is adiabatically equivalent to two conventional quantum spin Hall edges. Second is a strongly-correlated edge with only one pair of helical modes. The strongly-correlated edge also has a gap to single-electrons, but is gapless to pairs of electrons. In a quantum point contact geometry, this single-electron gap leads to a Fano factor of $2$ in shot noise measurements, compared to a Fano factor of $1$ for a weakly-correlated edge.
@@
Quantum spin Hall insulators (QSHIs) [1, 2, 3, 4, 5, 6] are prototypical examples of symmetry-protected topological phases in two dimensions. The edges of QSHIs characteristically hosts one-dimensional helical modes [7, 8, 9], which are guaranteed to be gapless as long as time-reversal ($\mathsf{T}$-) and/or $s^{z}$-spin conservation symmetry is preserved. If only $\mathsf{T}$-symmetry is present, QSHIs have a $\mathbb{Z}_{2}$ topological index [10, 11, 12], such that two copies of a QSHI form a trivial insulator. However, QSHIs that conserve the $s^{z}$-component of spin instead have a $\mathbb{Z}$ topological index [13], which is simply the $s^{z}$-weighted Chern number. In this case, two copies of a QSHI form a topologically non-trivial double quantum spin Hall insulator (DQSHI).
@@
Recently, there has been experimental evidence of DQSHIs in twisted homobilayers of the transition metal dichalcogenides (TMDs) WSe2 [14, 15] and MoTe2 [16, 17, 18]. The TMD-based DQSHIs can be understood as band insulators, where one fills the two lowest bands of the $K$ valley (which both have spin-up and Chern number $C=+1$) and the two lowest bands of the $K^{\prime}$ valley (which both have spin-down and $C=-1$). Since electronic interactions are not necessary for forming the bulk state, it is tempting to conclude that interactions have no effect at all. If this were true, TMD-based DQSHI edges would always have two pairs of helical edge modes, one per QSHI copy. However, twisted TMDs are strongly-correlated systems, and gapless topological edge modes can be heavily modified by interactions even when the bulk is stable [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. One therefore needs to carefully consider the effects of interactions in order to fully characterizing the edges of TMD-based DQSHIs.
strongly-correlated edge
weakly-correlated edge
double quantum spin Hall edge
edge interaction
irrelevant
relevant
edge interaction
@@
Figure 1: Effect of the interactions on the DQSHI edge. Without interactions, there are two pairs of helical edge modes (blue arrows indicate the direction of propagation, and gray arrows indicate spin of the mode). If the edge interactions (depicted in the dashed circle) are irrelevant, the number of helical modes is unchanged. If the interactions are relevant, the edge enters a strongly-correlated phase with only a single pair of helical modes that carry twice the charge and twice the spin of the original helical modes.
@@
In this work, we show that DQSHI edges have two symmetry-preserving phases if we include interactions: a weakly-correlated phase and a strongly-correlated phase. In the weakly-correlated phase, the edge has two pairs of helical modes, and is adiabatically connected to the non-interacting limit. The strongly-correlated phase, however, only has a single pair of helical modes. This edge also has a gap to adding or removing single electrons, but is gapless to adding or removing pairs of electrons. The phase transition between the weakly- and strongly-correlated phases occurs when a certain “double spin-flip” interaction term becomes relevant, as shown schematically in Fig. 1.
@@
Given that the DQSHI edge can realize two distinct phases, the next immediate question to ask is how one can distinguish them in an experimental setting. This is not possible via simple electrical transport, as the strongly- and weakly-correlated edges have the same quantized edge conductance. It is, however, possible to distinguish between the two phases using shot-noise measurements [31]. In a quantum point contact (QPC) geometry, the Fano factor of the current noise is $\mathcal{F}=1$ for the weakly-correlated edge (same as for a conventional QSHI or integer quantum Hall state [32, 33, 34, 35]). However, for the strongly-correlated edge, we instead find a doubled Fano factor of $\mathcal{F}=2$. This doubled Fano-factor cannot occur in non-interacting systems of electrons and is a clear indicator of strong edge correlations. A similar strongly correlated edge and doubled Fano-factor was also predicted to occur in bilayer graphene in a strong magnetic field[36, 37]. Here, due to the strong spin-orbit coupling of TMDs, we do not require an external magnetic field.
@@
Although we formally establish these results using the out-of-equilibrium Schwinger-Keldysh formalism, the doubled Fano factor of the strongly-correlated edge can be informally understood from two simple observations. First, due to the single-electron gap of the strongly-correlated edge, charge can only tunnel across the QPC in multiples of $2e$, as shown in Fig. 2. Second, the Fano factor is equal to the smallest amount of charge (in units of $e$) that can tunnel across the QPC [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. While the main focus of our work is on the double quantum spin Hall insulators realized in twisted TMDs, it is straightforward to generalize our arguments to other $N$-fold QSHIs, such as the triple ($N=3$) QSHIs, which has been observed in twisted MoTe2 [16, 18].
@@
This paper is organized as follows. In Section II, we introduce and bosonize the non-interacting DQSHI edge. In Section III we consider the effects of interactions on the DQSHI edge and show that, if certain interaction terms are relevant, the edge enters a strongly-correlated phase with a single-electron gap. In Section IV, we calculate the shot noise for the two possible DQSHI edges. We conclude our work in Section V, and discuss relevance to twisted TMD systems.
II Double quantum spin Hall edge.
@@
where $\bm{x}=(t,x)$ is the space-time coordinate, $a=1,2$ labels the two pairs of helical modes, and $\chi=\text{R$\uparrow$, L$\downarrow$}$ refers to the chirality/spin of the mode, and $k_{a\chi}$ is the Fermi-wavevector of the mode.
@@
$\mathcal{L}=-\frac{1}{4\pi}\partial_{x}\bm{\mathbf{\varphi}}^{\intercal}\bm{\mathbf{K}}\partial_{t}\bm{\mathbf{\varphi}}+\frac{1}{4\pi}\partial_{x}\bm{\mathbf{\varphi}}^{\intercal}\bm{\mathbf{V}}\partial_{x}\bm{\mathbf{\varphi}},$ (2)
where $\bm{\mathbf{\varphi}}$ is
$\bm{\mathbf{\varphi}}\equiv\begin{pmatrix}\varphi_{\text{1R$\uparrow$}}\\ \varphi_{\text{1L$\downarrow$}}\\ \varphi_{\text{2R$\uparrow$}}\\ \varphi_{\text{2L$\downarrow$}}\end{pmatrix},$ (3)
@@
$\bm{\mathbf{K}}\equiv\begin{pmatrix}1&&&\\ &-1&&\\ &&1&\\ &&&-1\end{pmatrix}.$ (4)
The matrix $\bm{\mathbf{V}}$ is the velocity matrix. For a non-interacting system, $\bm{\mathbf{V}}$ is a diagonal matrix with positive entries on the diagonal. As we shall show in the next section, $\bm{\mathbf{V}}$ generically has both diagonal and off-diagonal elements in interacting systems. Here, and throughout, we set $e=\hbar=1$, unless needed for clarity.
The DQSHI is guaranteed to have a gapless edge as long as $\mathrm{U}(1)_{c}$-charge conservation symmetry and $s^{z}$-spin conservation symmetry are preserved. Charge conservation symmetry acts on the bosonized theory as
@@
where $\theta_{s}$ is a spin dependent phase.
In addition to the $\mathrm{U}(1)_{c}$ and $s^{z}$ symmetries, we will also consider time-reversal symmetry, $\mathsf{T}$, which acts on the bosons as
@@
and sends $i$ to $-i$. It should be noted that while $\mathsf{T}$ is present in TMD DQSHIs, it is not a protecting symmetry, and the edge will be gapless even if $\mathsf{T}$ is broken, provided that the $\mathrm{U}(1)_{c}$- and $s^{z}$-conservations are preserved.
weakly-correlated
strongly-correlated
$e$
$2e$
Figure 2: Depiction of the quantum point contact geometry. In the weakly-correlated edge, the lowest possible charge that may tunnel across the point contact is $e$. However, for the strongly-correlated edge the minimal charge that can tunnel across the point contact is $2e$.
III The strongly-correlated edge.
@@
where we have suppressed the dependence of the bosons on the space-time coordinate $\bm{x}$. Since these density-density terms are quadratic in the bosonic fields, we can absorb them into the redefinition of $\bm{\mathbf{V}}$. Here we will consider a velocity matrix of the form
$\bm{\mathbf{V}}\rightarrow\begin{pmatrix}v&-v^{\prime}&u&-u^{\prime}\\ -v^{\prime}&v&-u^{\prime}&u\\ u&-u^{\prime}&v&-v^{\prime}\\ -u^{\prime}&u&-v^{\prime}&v\end{pmatrix},$ (9)
where $v$ is the renormalized velocity of the edge modes, $v^{\prime}$, $u$, $u^{\prime}$ correspond to various density interactions. This is not the most general velocity matrix, as we have explicitly omitted certain symmetry allowed density-density interactions. However, this form of $\bm{\mathbf{V}}$ will greatly simplify our forthcoming analysis. We provide an analysis of a general symmetric $\bm{\mathbf{V}}$ in Appendix C.
@@
As we shall show, if this interaction is relevant, the DQSHI edge will enter a non-trivial strongly-correlated phase. If the interaction is irrelevant then the DQSHI edge will remain adiabatically connected to the fully non-interacting limit.
@@
The values of $g_{-}$ and $v_{-}$ may also be renormalized, but this effect is not important for our present analysis. The values of $g_{+}$ and $v_{+}$ do not renormalize, as the channel-symmetric sector is a free bosonic theory, which is an RG fixed point. We will refer to this edge as the “weakly-correlated edge.”
@@
which consists of a single pair of bosons. We will refer to this as the “strongly-correlated edge”. Again, the values of $g_{+}$ and $v_{+}$ do not renormalize. It is worth explicitly noting that the $\varphi_{-}$ and $\vartheta_{-}$ bosons, do not carry $\mathrm{U}(1)_{c}$-charge or $s^{z}$-spin (see Eqs. 5,6 and 11), and so we have not broken any symmetries by gapping them out.
@@
In general, $k_{1\text{R}\uparrow}\neq k_{2\text{R}\uparrow}$ without fine-tuning. Because of this, the argument of the cosine term oscillates in space, and the interaction is RG-irrelevant. However, if $k_{1\text{R}\uparrow}=k_{2\text{R}\uparrow}$, $\mathcal{O}^{\prime}_{\text{dsf}}$ can be relevant, depending on the Luttinger parameter. Since the terms in Eq. 18 and Eq. 10 do not commute with each other, both terms cannot reach strong coupling simultaneously. Rather, there are two distinct strong coupling fixed points. The first fixed point is the one discussed above, where Eq. 10 flows to strong coupling, and Eq. 18 flows to weak coupling. In the second fixed point, Eq. 18 flows to strong coupling, and Eq. 10 flows to weak coupling. However, these two fixed points are equivalent under exchanging $\psi_{1\text{L}\downarrow}\leftrightarrow\psi_{2\text{L}\downarrow}$, and so we will not need to consider Eq. 18 in detail.
@@
An important feature of the strongly-correlated edge is that, despite being gapless, it is gapped to single electron operators. By contrast, the weakly-correlated edge is gapless to single electrons, as in the decoupled limit.
@@
For the weakly-correlated edge, all bosons are free, and so $\psi^{\dagger}_{1\text{R}\uparrow}$ has power-law correlations [48], indicating that it is gapless. However, for the strongly-correlated edge, $\vartheta_{-}$ is pinned to its minimum, and $\varphi_{-}$ has short-ranged correlations. Therefore, $\psi^{\dagger}_{1\text{R}\uparrow}$ has short-ranged electronic correlations in the strongly-correlated edge. By the same reasoning, the three other electron operators from Eq. 1 also have short-ranged correlations. Individual electrons are therefore gapped in the strongly-correlated regime.
@@
where we have used that $\vartheta_{-}$ acquires a classical expectation value in the strongly-correlated phase ($\vartheta_{-}=\langle\vartheta_{-}\rangle$). By inspection, the operators in Eq. 21 are bosons with charge $2e$ and spin ±$\pm 1$, the operator in Eq. 22 is a boson with charge $2e$ and spin 00, and the operator in Eq. 23 is a boson with charge 00 and spin $1$. Since the operators are all independent of $\varphi_{-}$, they are all gapless with power law correlations.
Furthermore, all other local operators in the strongly-correlated edge can be written as combinations of the bosons in Eqs. 21-23. This can be directly argued from the fact that any gapless degree of freedom must be independent of $\varphi_{-}$, and that any local operator (i.e., one made out of a finite number of electron operators) must be invariant under the transformation $\varphi_{a\chi}\rightarrow\varphi_{a\chi}+2\pi$, for all $a$ and $\chi$ individually.
IV Distinguishing the strongly- and weakly-correlated edges via shot noise.
We now ask how one can differentiate between the two possible edge phases of a TMD DQSHI. Using the Kubo formula, one finds that both the weakly- and strongly-correlated edges have edge conductance $2e^{2}/h$, (assuming the leads are Fermi liquid–like) [51]. However, as we shall show in this section, the edges can be distinguished by the current shot noise in a quantum point contact (QPC) geometry. In the QPC geometry, the edge states are brought together at a constriction which separates the source from the drain, allowing for tunneling between the top and bottom edges (see Fig. 2). The strength of tunneling can be tuned by via a gate voltage [52, 53]. Here, we will assume the QPC is in the weak backscattering limit [38]. The analysis in the weak tunneling limit is nearly identical, as the bulk of the DQSHI is not topologically ordered. Specifically, the Fano factor is the same in the weak backscattering and weak tunneling limits.
@@
where $\mathcal{F}$ is the Fano factor. For topological systems dominated by edge transport, $\mathcal{F}$ is expected to be equal to the smallest charge (in units of $e$) that can tunnel across the QPC at low voltages. Since the strongly-correlated edge is gapped to all individual single-electrons, charge must necessarily be backscattered in units of $2e$. For the weakly-correlated edge, there are gapless electrons and charges can be backscattered in units of $e$. Based on this, we expect that $\mathcal{F}=2$ for the strongly-correlated edge, and $\mathcal{F}=1$ for the weakly-correlated edge. In the following, we will explicitly show that this is indeed the case, and that shot noise is a robust, experimentally-accessible signature that can distinguish between the two possible DQSHI edges.
@@
$\begin{split}\varphi_{\text{c}}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow+}-\varphi_{\text{L}\uparrow+})+(\varphi_{\text{R}\downarrow+}-\varphi_{\text{L}\downarrow+})\right]\\ \vartheta_{\text{c}}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow+}+\varphi_{\text{L}\uparrow+})+(\varphi_{\text{R}\downarrow+}+\varphi_{\text{L}\downarrow+})\right]\\ \varphi_{\text{s}}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow+}-\varphi_{\text{L}\uparrow+})-(\varphi_{\text{R}\downarrow+}-\varphi_{\text{L}\downarrow+})\right]\\ \vartheta_{\text{s}}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow+}+\varphi_{\text{L}\uparrow+})-(\varphi_{\text{R}\downarrow+}+\varphi_{\text{L}\downarrow+})\right]\\ \varphi_{\text{n}\pm}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow-}-\varphi_{\text{L}\uparrow-})\pm(\varphi_{\text{R}\downarrow-}-\varphi_{\text{L}\downarrow-})\right]\\ \vartheta_{\text{n}\pm}&\equiv\frac{1}{2}\left[(\varphi_{\text{R}\uparrow-}+\varphi_{\text{L}\uparrow-})\pm(\varphi_{\text{R}\downarrow-}+\varphi_{\text{L}\downarrow-})\right]\\ \end{split}$ (27)
@@
In terms of the new bosonic fields, the Lagrangian density for the weakly-correlated edge is:
@@
$\begin{split}\mathcal{L}_{0,\text{cs}}=&\frac{v_{\text{c}}}{4\pi}\left[g_{\text{c}}(\partial_{x}\vartheta_{\text{c}})^{2}+\frac{1}{g_{\text{c}}}(\partial_{x}\varphi_{\text{c}})^{2}\right]\\ +&\frac{v_{\text{s}}}{4\pi}\left[g_{\text{s}}(\partial_{x}\vartheta_{\text{s}})^{2}+\frac{1}{g_{\text{s}}}(\partial_{x}\varphi_{\text{s}})^{2}\right]\\ \mathcal{L}_{0,\text{n}}=&\frac{v_{\text{n}+}}{4\pi}\left[g_{\text{n}+}(\partial_{x}\vartheta_{\text{n}+})^{2}+\frac{1}{g_{\text{n}+}}(\partial_{x}\varphi_{\text{n}+})^{2}\right]\\ +&\frac{v_{\text{n}-}}{4\pi}\left[g_{\text{n}-}(\partial_{x}\vartheta_{\text{n}-})^{2}+\frac{1}{g_{\text{n}-}}(\partial_{x}\varphi_{\text{n}-})^{2}\right]\\ \end{split}$ (29)
@@
$\begin{split}g_{\text{c}}=g_{+}=1/g_{\text{s}},\\ g_{\text{n}+}=g_{-}=1/g_{\text{n}-}.\\ \end{split}$ (30)
@@
[…]
The neutral sector is fully gapped for the strongly-correlated edge, and so the low energy Lagrangian for the strongly-correlated edge is simply $\mathcal{L}_{0,\text{cs}}$.

### v3

# Shot noise in strongly correlated double quantum spin Hall edges

@@ Abstract
We consider the effects of interactions on the edges of “double” quantum spin Hall insulators (DQSHIs), motivated by recent experiments on moiré twisted metal dichalcogenides. Without interactions, a DQSHI can be understood as two copies of a conventional quantum spin Hall insulator. If interactions are present and $s^{z}$-spin is conserved, we show that there are two possible phases for the DQSHI edge. First is a weakly correlated edge which has two pairs of helical modes and is adiabatically equivalent to two conventional quantum spin Hall edges. Second is a strongly correlated edge with only one pair of helical modes. The strongly correlated edge also has a gap to single electrons, but is gapless to pairs of electrons. In a quantum point contact geometry, this single-electron gap leads to a Fano factor of $2$ in shot noise measurements, compared to a Fano factor of $1$ for a weakly correlated edge.
@@
Quantum spin Hall insulators (QSHIs) [1, 2, 3, 4, 5, 6] are prototypical examples of symmetry-protected topological phases in two dimensions. The edges of QSHIs characteristically host one-dimensional helical modes [7, 8, 9], which are guaranteed to be gapless as long as time-reversal symmetry ($\mathsf{T}$-symmetry) is present and/or the $s^{z}$-component of spin is conserved. If only $\mathsf{T}$-symmetry is present, QSHIs have a $\mathbb{Z}_{2}$ topological index [10, 11, 12], such that two copies of a QSHI form a trivial insulator. However, QSHIs that conserve the $s^{z}$-component of spin instead have a $\mathbb{Z}$ topological index [13], which is simply the $s^{z}$-weighted Chern number. In this case, two copies of a QSHI form a topologically non-trivial double quantum spin Hall insulator (DQSHI) [14, 15, 16].
@@
Recently, there has been experimental evidence of DQSHIs in twisted homobilayers of the transition metal dichalcogenides (TMDs) WSe2 [17, 18] and MoTe2 [19, 20, 21]. The TMD-based DQSHIs can be understood as band insulators, where one fills the two lowest bands of the $K$ valley (which both have spin-up and Chern number $C=+1$) and the two lowest bands of the $K^{\prime}$ valley (which both have spin-down and C=1C=-1). Since electronic interactions are not necessary for forming the bulk state, it is tempting to conclude that interactions have no effect at all. If this were true, TMD-based DQSHI edges would always have two pairs of helical edge modes, one per QSHI copy. However, twisted TMDs are strongly correlated systems, and gapless topological edge modes can be heavily modified by interactions even when the bulk is stable [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. One therefore needs to carefully consider the effects of interactions in order to fully characterize the edges of TMD-based DQSHIs.
@@
Figure 1: (a) Effect of the interactions on the DQSHI edge. Without interactions, there are two pairs of helical edge modes (blue arrows indicate the direction of propagation, and gray arrows indicate spin of the mode). If the edge interactions (depicted in the dashed circle) are irrelevant, the number of helical modes is unchanged. If the interactions are relevant, the edge enters a strongly correlated phase with only a single pair of helical modes that carry twice the charge and twice the spin of the original helical modes. (b) Edge interaction depicted in energy-momentum space.
@@
In this work, we show that DQSHI edges have two symmetry-preserving phases if we include interactions: a weakly correlated phase and a strongly correlated phase. In the weakly correlated phase, the edge has two pairs of helical modes, and is adiabatically connected to the non-interacting limit. The strongly correlated phase, however, only has a single pair of helical modes. This edge also has a gap to adding or removing single electrons, but is gapless to adding or removing pairs of electrons. The phase transition between the weakly and strongly correlated phases occurs when a certain “double spin-flip” interaction term becomes relevant, as shown schematically in Fig. 1.
@@
Given that the DQSHI edge can realize two distinct phases, the next immediate question to ask is how one can distinguish them in an experimental setting. This is not possible via simple electrical transport, as the strongly and weakly correlated edges have the same quantized edge conductance. It is, however, possible to distinguish between the two phases using shot noise measurements [37, 38, 39, 40, 41]. In a quantum point contact (QPC) geometry, the Fano factor of the current noise is $\mathcal{F}=1$ for the weakly correlated edge (same as for a conventional QSHI or integer quantum Hall state [42, 43, 44, 45]). However, for the strongly correlated edge, we instead find a doubled Fano factor of $\mathcal{F}=2$. This doubled Fano factor cannot occur in non-interacting systems of electrons and is a clear indicator of strong edge correlations. A similar strongly correlated edge and doubled Fano-factor was also predicted to occur in bilayer graphene in a strong magnetic field [46, 47]. We do not require an external magnetic field, since TMDs have strong Ising spin-orbit coupling [48, 49, 50].
@@
Although we formally establish these results using the out-of-equilibrium Schwinger-Keldysh formalism, the doubled Fano factor of the strongly correlated edge can be informally understood from two simple observations. First, due to the single-electron gap of the strongly correlated edge, charge can only tunnel across the QPC in multiples of $2e$, as shown in Fig. 2. Second, the Fano factor is equal to the smallest amount of charge (in units of $e$) that can tunnel across the QPC [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. While the main focus of our work is on the double quantum spin Hall insulators realized in twisted TMDs, it is straightforward to generalize our arguments to other $N$-fold QSHIs, such as the triple ($N=3$) QSHIs, which has been observed in twisted MoTe2 [19, 21].
@@
This paper is organized as follows. In Section II, we introduce and bosonize the non-interacting DQSHI edge. In Section III we consider the effects of interactions on the DQSHI edge and show that, if certain interaction terms are relevant, the edge enters a strongly correlated phase with a single-electron gap. In Section IV, we calculate the shot noise for the two possible DQSHI edges. We conclude our work in Section V, and discuss relevance to twisted TMD systems.
II Bosonization of the double quantum spin Hall edge.
@@
where $\bm{x}=(t,x)$ is the space-time coordinate, $a=1,2$ labels the two pairs of helical modes, and $\chi=\text{R$\uparrow$, L$\downarrow$}=+1,-1$ refers to the chirality/spin of the mode, and $k_{a\chi}$ is the Fermi-wavevector of the mode.
@@
$\mathcal{L}=-\frac{1}{4\pi}\partial_{x}\boldsymbol{\mathbf{\varphi}}^{\intercal}\boldsymbol{\mathbf{K}}\partial_{t}\boldsymbol{\mathbf{\varphi}}+\frac{1}{4\pi}\partial_{x}\boldsymbol{\mathbf{\varphi}}^{\intercal}\boldsymbol{\mathbf{V}}\partial_{x}\boldsymbol{\mathbf{\varphi}},$ (2)
where $\boldsymbol{\mathbf{\varphi}}$ is
$\boldsymbol{\mathbf{\varphi}}\equiv\begin{pmatrix}\varphi_{\text{1R$\uparrow$}}\\ \varphi_{\text{1L$\downarrow$}}\\ \varphi_{\text{2R$\uparrow$}}\\ \varphi_{\text{2L$\downarrow$}}\end{pmatrix}$, (3)
@@
$\boldsymbol{\mathbf{K}}\equiv\begin{pmatrix}1&&&\\ &-1&&\\ &&1&\\ &&&-1\end{pmatrix}$. (4)
The matrix $\boldsymbol{\mathbf{V}}$ is the velocity matrix. For a non-interacting system, $\boldsymbol{\mathbf{V}}$ is a diagonal matrix with positive entries on the diagonal. As we shall show in the next section, $\boldsymbol{\mathbf{V}}$ generically has both diagonal and off-diagonal elements in interacting systems. Here, and throughout, we set $e=\hbar=1$, unless needed for clarity.
The DQSHI is guaranteed to have a gapless edge as long as charge and the $s^{z}$-component of spin are conserved. In twisted TMDs, first-principles calculations show that $s^{z}$ is indeed a good quantum number, since the spin-orbit coupling is dominantly Ising-like [48, 49, 50, 63]. First-principles calculations also indicate that $s^{z}$ is a good quantum number even when the out of plane mirror symmetry is broken by an external displacement field[64]. The conservation of charge and the $s^{z}$-component of spin defines two $\mathrm{U}(1)$-symmetries, which we will refer to as $\mathrm{U}(1)_{\text{c}}$-charge conservation symmetry and $s^{z}$-spin conservation symmetry, respectively [65]. The charge-conservation symmetry acts on the fermion operators as
$\psi_{a\chi}\to e^{i\theta_{c}}\psi_{a\chi},$ (5)
whereas spin-conservation symmetry acts axially on the different fermion species,
$\psi_{a\chi}\to e^{i\chi\theta_{s}}\psi_{a\chi}.$ (6)
Using the correspondence in equation (1), one finds that the symmetry operators act on the bosonic modes as
@@
where $\theta_{s}$ is a spin-dependent phase.
In addition to the $\mathrm{U}(1)_{c}$ and $s^{z}$ symmetries, we will also consider time-reversal symmetry, $\mathsf{T}$. This acts on the fermionic operators as
$\psi_{a\chi}\to\chi\psi_{a\bar{\chi}}$ (9)
and sends $i$ to i-i, where $\bar{\chi}$ stands for the opposite chirality/spin label. Hence, the bosons transform as
@@
It should be noted that while $\mathsf{T}$ is present in TMD DQSHIs, it is not a protecting symmetry, and the edge will be gapless even if $\mathsf{T}$ is broken, provided that the $\mathrm{U}(1)_{c}$- and $s^{z}$-conservations are preserved.
Figure 2: Depiction of the quantum point contact geometry. In the weakly correlated edge, the lowest possible charge that may tunnel across the point contact is $e$, via any of the terms in Eq. 36. However, for the strongly correlated edge the minimal charge that can tunnel across the point contact is $2e$.
III The strongly correlated edge.
@@
where we have suppressed the dependence of the bosons on the space-time coordinate $\bm{x}$. Since these density-density terms are quadratic in the bosonic fields, we can absorb them into the redefinition of $\boldsymbol{\mathbf{V}}$. Here we will consider a velocity matrix of the form
$\boldsymbol{\mathbf{V}}\rightarrow\begin{pmatrix}v&-v^{\prime}&u&-u^{\prime}\\ -v^{\prime}&v&-u^{\prime}&u\\ u&-u^{\prime}&v&-v^{\prime}\\ -u^{\prime}&u&-v^{\prime}&v\end{pmatrix}$, (12)
where $v$ is the renormalized velocity of the edge modes, $v^{\prime}$, $u$, $u^{\prime}$ correspond to various density interactions. This is not the most general velocity matrix, as we have explicitly omitted certain symmetry allowed density-density interactions. However, this form of $\boldsymbol{\mathbf{V}}$ will greatly simplify our forthcoming analysis. We provide an analysis of a general symmetric $\boldsymbol{\mathbf{V}}$ in Appendix C.
@@
As we shall show, if this interaction is relevant, the DQSHI edge will enter a non-trivial strongly correlated phase. If the interaction is irrelevant then the DQSHI edge will remain adiabatically connected to the fully non-interacting limit.
@@
The values of $g_{-}$ and $v_{-}$ may also be renormalized, but this effect is not important for our present analysis. The values of $g_{+}$ and $v_{+}$ do not renormalize, as the channel-symmetric sector is a free bosonic theory, which is an RG fixed point. We will refer to this edge as the “weakly correlated edge.”
@@
which consists of a single pair of bosons. We will refer to this as the “strongly correlated edge”. Again, the values of $g_{+}$ and $v_{+}$ do not renormalize. It is worth explicitly noting that the $\varphi_{-}$ and $\vartheta_{-}$ bosons, do not carry $\mathrm{U}(1)_{c}$-charge or $s^{z}$-spin (see Eqs. 7,8 and 14), and so we have not broken any symmetries by gapping them out.
@@
In general, $k_{1\text{R}\uparrow}\neq k_{2\text{R}\uparrow}$ without fine-tuning. Because of this, the argument of the cosine term oscillates in space, and the interaction is RG-irrelevant. However, if $k_{1\text{R}\uparrow}=k_{2\text{R}\uparrow}$, $\mathcal{O}^{\prime}_{\text{dsf}}$ can be relevant, depending on the Luttinger parameter. Since the terms in Eq. 21 and Eq. 13 do not commute with each other, both terms cannot reach strong coupling simultaneously. Rather, there are two distinct strong coupling fixed points. The first fixed point is the one discussed above, where Eq. 13 flows to strong coupling, and Eq. 21 flows to weak coupling. In the second fixed point, Eq. 21 flows to strong coupling, and Eq. 13 flows to weak coupling. However, these two fixed points are equivalent under exchanging $\psi_{1\text{L}\downarrow}\leftrightarrow\psi_{2\text{L}\downarrow}$. Since one can arbitrarily assign labels to the two channels in the fine-tuned scenario, we will not need to consider Eq. 21 in detail.
@@
An important feature of the strongly correlated edge is that, despite being gapless, it is gapped to single electron operators. By contrast, the weakly correlated edge is gapless to single electrons, as in the decoupled limit.
@@
For the weakly correlated edge, all bosons are free, and so $\psi^{\dagger}_{1\text{R}\uparrow}$ has power-law correlations [61], indicating that it is gapless. However, for the strongly correlated edge, $\vartheta_{-}$ is pinned to its minimum, and $\varphi_{-}$ has short-ranged correlations. Therefore, $\psi^{\dagger}_{1\text{R}\uparrow}$ has short-ranged electronic correlations in the strongly correlated edge. By the same reasoning, the three other electron operators from Eq. 1 also have short-ranged correlations. Individual electrons are therefore gapped in the strongly correlated regime.
@@
where we have used that $\vartheta_{-}$ acquires a classical expectation value in the strongly correlated phase ($\vartheta_{-}=\langle\vartheta_{-}\rangle$). By inspection, the operators in Eq. 24 are bosons with charge $2e$ and spin $\pm 1$, the operator in Eq. 25 is a boson with charge $2e$ and spin 0, and the operator in Eq. 26 is a boson with charge 0 and spin $1$. Since the operators are all independent of $\varphi_{-}$, they are all gapless with power law correlations.
Furthermore, all other local operators in the strongly correlated edge can be written as combinations of the bosons in Eqs. 24-26. This can be directly argued from the fact that any gapless degree of freedom must be independent of $\varphi_{-}$, and that any local operator (i.e., one made out of a finite number of electron operators) must be invariant under the transformation $\varphi_{a\chi}\rightarrow\varphi_{a\chi}+2\pi$, for all $a$ and $\chi$ individually.
IV Distinguishing the strongly and weakly correlated edges via shot noise.
We now ask how one can differentiate between the two possible edge phases of a TMD DQSHI. Using the Kubo formula, one finds that both the weakly and strongly correlated edges have edge conductance $2e^{2}/h$, (assuming the leads are Fermi liquid–like) [67]. However, as we shall show in this section, the edges can be distinguished by the current shot noise in a quantum point contact (QPC) geometry. In the QPC geometry, the edge states are brought together at a constriction which separates the source from the drain, allowing for tunneling between the top and bottom edges (see Fig. 2). The strength of tunneling can be tuned by via a gate voltage [68, 69]. Here, we will assume the QPC is in the weak backscattering limit [51, 70]. The analysis in the weak tunneling limit is nearly identical, as the bulk of the DQSHI is not topologically ordered. Specifically, the Fano factor is the same in the weak backscattering and weak tunneling limits.
@@
where $\mathcal{F}$ is the Fano factor. For topological systems dominated by edge transport, $\mathcal{F}$ is expected to be equal to the smallest charge (in units of $e$) that can tunnel across the QPC at low voltages. Since the strongly correlated edge is gapped to all individual single electrons, charge must necessarily be backscattered in units of $2e$. For the weakly correlated edge, there are gapless electrons and charges can be backscattered in units of $e$. Based on this, we expect that $\mathcal{F}=2$ for the strongly correlated edge, and $\mathcal{F}=1$ for the weakly correlated edge. In the following, we will explicitly show that this is indeed the case, and that shot noise is a robust, experimentally-accessible signature that can distinguish between the two possible DQSHI edges.
@@
[…]
@@
In terms of the new bosonic fields, the Lagrangian density for the weakly correlated edge is:
@@
[…]
@@
[…]
@@
with $\gamma=\text{c, s}$. The Lagrangian for the neutral sector, $\mathcal{L}_{0,\text{n}}$, is defined analogously.
The neutral sector is fully gapped for the strongly correlated edge, and so the low energy Lagrangian for the strongly correlated edge is simply $\mathcal{L}_{0,\text{cs}}$.